OpenStudy (anonymous):

What is the equation of the line perpendicular to the line y =4 over 9 x – 2 that passes through the point (4,3)?

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

o u want me to give u the slop ok 1 sec 4/9

OpenStudy (anonymous):

how do i get that one

OpenStudy (anonymous):

ok so wont the anser be y=-4/9x-2

OpenStudy (anonymous):

this is a mult qwestion it is one of thie right y = –x – 2 y = -4/9x - 2 y = 9/4x + 1 y = –9/4x – 12 y = –9/4x +12

OpenStudy (anonymous):

ok ty